OpenStudy (anonymous):

how do you factor x^2+5x-14=0

hero (hero):

I guess you had trouble understanding the approach I showed you

OpenStudy (anonymous):

i cant factor to save my life sir.... and no mater what i try i get it wrong every time.... :(

hero (hero):

Let me post the solution, then you can ask me questions about what you don't get.

OpenStudy (anonymous):

ok. thank you.

OpenStudy (anonymous):

ok i understand it... but i always have to have help for some reason.... i can do perfect square ones but not the ones that i am doing now. and it sucks....

OpenStudy (anonymous):

hey hero... acording to the thing it is -7,2 and not the other one.

hero (hero):

x^2+5x-14=0 x^2-2x+7x-14 = 0 x(x-2)+7(x-2) = 0 (x-2)(x+7) = 0 x = -7,2

hero (hero):

It's a good thing you have a way to check your answers

OpenStudy (anonymous):

ya. :)

hero (hero):

If you can't figure any of the other ones out, then I'm afraid that you don't fully understand the process.

OpenStudy (anonymous):

unfortuantly..... i hate factoring...